OpenStudy (poofypenguin):

How many vertical asymptotes does the function f(x)=ln(|(x+1)/(x^2-7x+12)| have?

OpenStudy (campbell_st):

let the denominator equal zero, then solve for x \[x^2 - 7x + 12 = 0\] this can be done by factorising the solutions will be the vertical asymptotes

OpenStudy (poofypenguin):

Alright...I understand that vertical asymptotes come from factoring the denominator, but the thing that's confusing me is that my solution manual says that the numerator will also cause a vertical asymptote... is that possible? Could you explain that to me, please?

OpenStudy (anonymous):

Numerator equaling 0 would cause a horizontal asymptote I believe.

OpenStudy (poofypenguin):

Are you sure? Cuz both my prof and my answer key says that it'll be a vertical asymptote...I'm so confused right now...:/

OpenStudy (campbell_st):

the numerator may cause a point of discontinuity.. e.g. \[\frac{(x + 1)}{(x + 1)(x + 5)}\] (x + 1) is a common factor... so the curve is discontinuous at that point. a horizontal asymptote may come form the numerator... and is easily found by division.

OpenStudy (campbell_st):

oops I missed that it was a log function and with a log function you can also get a asymptote from the numerator... x = -1 they are the values tha make it undefined... x = 3, and x = 4 and given its a log function... you can't find the log of zero. so the numerator can't be zero so solve x + 1 = 0 to find the asymptote. x = -1 so x can't be -1 hope that helps

OpenStudy (tkhunny):

It's a logarithm! What happens when the argument of the logarithm is zero (0)?

OpenStudy (campbell_st):

hope that makes some sense

OpenStudy (poofypenguin):

Yes! I get it now! Thanks so much! :D